currency

Spot XR

Change ( from day previous)

6 mos. Fwd

CAD (per $US)

1.3155

-0.0032

1.3112

AUD (in $US)

0.7267

+0.0036

0.7287

Jap. yen (per $US)

110.53

-0.37

110.5147

See the exchange rate information above. Some currencies are quoted in USD, some per USD. The “change” is from the previous day.

  1. Indicate for each currency whether it appreciated or depreciated from the day previous.
  2. Calculate the change in the USD value of 15 million units of each currency above from the day previous.
  3. For each currency above, indicate whether the currency is trading at a forward premium or a forward discount.
Answer #1

a):- Indicate for each currency whether it appreciated or depreciated from the day previous.

The change (from day previous) with respect to CAD (per US$) is negative.

Negative change implies fewer CAD will be exchanged per US$ relative to day previous.

When fewer units of a domestic currency is exchanged for 1 unit of foreign currency relative to the previous time period then such domestic currency is said to be appreciating.

So, The CAD is appreciated from the day previous.

The change (from day previous) with respect to AUD (per US$) is positive.

Positive change implies more AUD will be exchanged per US$ relative to day previous.

When more units of a domestic currency is exchanged for 1 unit of foreign currency relative to the previous time period then such domestic currency is said to be depreciating.

So,The AUD is depreciated from the day previous.

The change (from day previous) with respect to Japanese Yen (per US$) is negative.

Negative change implies fewer Yens will be exchanged per US$ relative to day previous.

When fewer units of a domestic currency is exchanged for 1 unit of foreign currency relative to the previous time period then such domestic currency is said to be appreciating.

So,The Japanese Yen is appreciated from the day previous.

b):- Calculate the change in the USD value of 15 million units of each currency above from the day previous.

USD = 1.3155 CAD

CAD 15 million = USD 15 million / 1.3155

or, CAD 15 million = USD 11.40 million

New Excahnge Rate is 1 USD = 1.3155-0.0032 = 1.3123 CAD

New CAD 15 million = USD 15 million / 1.3123 =USD 11.43 million

Hence CAD 15 million which is USD 11.40 million will change to USD 11.43 million

1 USD = 0.7267 AUD

AUD 15 million = USD 15 million / 0.7267

or, AUD 15 million = USD 20.64 million

New Excahnge Rate is 1 USD = 0.7267+0.0036 = 0.7303 AUD

New AUD 15 million = USD 15 million / 0.7303 =USD 20.54 million

Hence AUD 15 million which is USD 20.64 million will change to USD 20.54 million

1 USD = 110.53 Jap. Yen

Jap Yen 15 million = USD 15 million / 110.53 = USD 0.1357 million

or, Jap Yen 15 million = USD 0.1357 million

New Excahnge Rate is 1 USD = 110.53-0.37= 110.16 Jap Yen

NewJap Yen 15 million = USD 15 million / 110.16 =USD 0.1362 million

Hence Jap Yen 15 million which is USD 0.1357 million will change to USD 0.1362 million

c):- For each currency above, indicate whether the currency is trading at a forward premium or a forward discount.

A currency is said to be trading at forward premium when forward exchange rate is greater than the spot exchange rate.

A currency is said to be trading at forward discount when forward exchange rate is less than the spot exchange rate.

The 6 Mos. Fwd exchange rate (1.3112 CAD per US$) of CAD is less than the spot exchange rate (1.3155 CAD per US$) of CAD.

So,The CAD is trading at a forward discount.

The 6 Mos. Fwd exchange rate (0.7287 AUD per US$) of AUD is more than the spot exchange rate (0.7267 AUD per US$) of AUD.

So,The AUD is trading at a forward premium.

The 6 Mos. Fwd exchange rate (110.5147 Japanese yen per US$) of Japanese yen is less than the spot exchange rate (110.53 Japanese yen per US$) Japanese yen.

So,The Japanese yen is trading at a forward discount.

please like the answer...